A 78 g container of containing a liquid solution is heated to 98⁰C. Another 3g of the same solution at temperature of 40⁰C is added to the container. Determine the final temperature of the solution?


1
Expert's answer
2022-01-18T09:35:52-0500

By thermal equilibrium:


"cm_1(t_1-t)+cm_2(t_2-t)=0,\\\space\\
t=\dfrac{m_1t_1+m_2t_2}{m_1+m_2}=96°\text C."